Find the coordinates of the midpoint of a segment with the given endpoints. Please show all work.

J(-1,7)

K(3,-3)

thank you

Answer by stanbon(29602)

Find the coordinates of the midpoint of a segment with the given endpoints. Please show all work.
J(-1,7)
K(3,-3)
-------------------------------
Let the midpoint coordinates be (x,y)
---
x = (-1+3)/2 = 1
y = (7-3)/2 = 2
-----------------------
midpoint is (1,2)
================================
Cheers,
Stan H.